OpenStudy (anonymous):

Will give medal and fan . One is made by dissolving lithium carbonate (Ksp = 8.15 x 10-4) in 100.0 mL of water until excess solid is present, while the other is prepared by dissolving lithium phosphate (Ksp = 2.37 x 10-4) in 200.0 mL of water until excess solid is present. Which solution has a higher lithium cation concentration?

OpenStudy (anonymous):

Li2CO3 8.15 * 10 ^ -4 = 4x^3 x = 0.0588 Li concentration : 0.1176 Li3PO4 2.37 * 10^-4 = (3x)^3 * x x = cube root of ((2.37*10^-4) / 9 ) x = 0.0298 Concentration of Lithum = 0.089 BUT It is wrong the concentration of Li in Phostphate should be more than the first , where am I wrong ??? PLZ TELL .

OpenStudy (anonymous):

@ganeshie8 plz help

OpenStudy (anonymous):

@anonymous_user @Squirrels @em0ney @thomaster @thadyoung @Destinycholie @Compassionate

OpenStudy (anonymous):

Using the method mentioned at : https://www.chem.tamu.edu/class/majors/tutorialnotefiles/ksp.htm

OpenStudy (anonymous):

@ganeshie8 @Abhisar @thomaster @Compassionate @radar @paki

OpenStudy (anonymous):

@ganeshie8 @thomaster @Auxuris @Compassionate @radar @ParthKohli @Preetha

OpenStudy (anonymous):

@DUDE..IM..A..DUCK @sammixboo @Secret-Ninja

OpenStudy (anonymous):

@Auxuris @alyssa_xo @ayee_ciera @aha.its.kristen @studygeek15 @stonewoods @sleepyjess @Secret-Ninja @Squirrels

OpenStudy (secret-ninja):

Sorry, I'm not in Chemistry! :/

OpenStudy (sleepyjess):

sorry I haven't learned this yet